A study of the time spent shopping in a supermarket for 20 specific items showed an approximately uniform distribution between 25 25 minutes and 45 45 minutes. Bold a. What a. What is the probability that the shopping time will be between 30 30 and 37 37 ​minutes? Bold b. What b. What is the probability that the shopping time will be less than 36 36 ​minutes? Bold c. What c. What are the mean and standard deviation of the shopping​ time? a. The probability that the shopping time will be between 30 30 and 37 37 minutes is nothing . ​(Type an integer or a​ decimal.) b. The probability that the shopping time will be less than 36 36 minutes is nothing . ​(Type an integer or a​ decimal.) c. The mean of the given uniform distribution is mu μ equals = nothing . ​(Type an integer or a​ decimal.) The standard deviation of the given uniform distribution is sigma σ equals = nothing . ​(Round to four decimal places as​ needed.)
Answer #1

a) P(Between 30 and 37) = (37- 30)/(45 - 25) = 7/20 = 0.35

b) P(Less than 36) = (36 - 25)/(45 - 25) = 11/20 = 0.55

c) Mean = (25 + 45)/2 = 35

Standard deviation = = 5.7735